OpenStudy (anonymous):

hey guys have you done all the Midterm ? for me is just missing Q4, Q5, Q6 ... Does anybody help me finish these ? Q4 : What is the expression for vOUT in terms of vIN, R, and b when vIN is greater than 0V? What is the expression for vOUT in terms of vIN, VS, R, and b for this amplifier when vIN is greater than 0V?

OpenStudy (anonymous):

first, (R*b)/(1+R*b)*vIN : second (VS*(1+R*b)-vIN)/(1+R*b)

OpenStudy (anonymous):

Thank so much Malchiva !! .. Can you help me with Q5 and Q6 too ?

OpenStudy (anonymous):

1)Vo* (R3-(B*R1*R3)+R1+R2+A*R2)/(R3-B*R1*R3)=VIN 2)i1=V0/(R3-B*R1*R3) 3)vo = vth 4)Rth=Vth/Isc

OpenStudy (anonymous):

Q6: The circuit elements have the following values: Vin=15V, R1=0.5Ω, R2=1.5Ω, R3=5Ω, A=2 and B=1.5. is it formula of Q6 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you very much Vivek... I'll solve this.. after I'll say the answers here.. after can you check this out whether the values correct or not ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Vo = 3 ?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

I= 2.4 ? vth= 3 Rth = 1.25 Ohm ?

OpenStudy (anonymous):

not sure of Rth but other answers are correct.. just give it a try

OpenStudy (anonymous):

yeah... you're right the Rth is not correct..

OpenStudy (anonymous):

Do you know how can I solve it ? and by the way.. you know how Can I solve the Q5 ?

OpenStudy (anonymous):

a)vs-(r*iD) b)-3*f*(Vgs^2)*R

OpenStudy (anonymous):

try Rth to be 2

OpenStudy (anonymous):

sorry, but it's no correct .. it's not 2

OpenStudy (anonymous):

i dont know:( got Q5?

OpenStudy (anonymous):

and I only have one more chance.. because it's my final check right now on Q6

OpenStudy (anonymous):

yeeah I got it .. thanks very much man

OpenStudy (anonymous):

your values on Q5 is the same of mine ?

OpenStudy (anonymous):

wat r ur values?

OpenStudy (anonymous):

Here, VS=80V, R=1.5kΩ, and the transistor parameter f=2mAV3. The amplifier is biased with VIN=2V.

OpenStudy (anonymous):

R is 3K

OpenStudy (anonymous):

Rth is it 3 ? on Q6 ? are you sure.. because its my final check on Q6 :/

OpenStudy (anonymous):

3k ?

OpenStudy (anonymous):

sry its my r value in Q5

OpenStudy (anonymous):

i'm not sure of Rth